OpenStudy (anonymous):

Graph the equation by plotting points y=1/2x-3 (2,_),(0,_). I think the answers are -2 and -3 but I am not positive about this. Could someone please let me know if this is correct?

OpenStudy (anonymous):

just fit it back in and see what you get

OpenStudy (anonymous):

the equation you have is y=1/2x-3 you have points these are only values for x (2,_) (0,_).

OpenStudy (anonymous):

x | y ------- | | | if you remembered working for something like this

OpenStudy (anonymous):

with*

OpenStudy (anonymous):

yes but I am not sure that I did it correctly, that is why I need someone to verify it for me to be sure that it is correct

jimthompson5910 (jim_thompson5910):

y = (1/2)x-3 y = (1/2)(2)-3 ... plug in x = 2 y = 2/2 - 3 y = 1 - 3 y = -2 So (2, -2) is one ordered pair

jimthompson5910 (jim_thompson5910):

y = (1/2)x-3 y = (1/2)(0)-3 ... plug in x = 0 y = 0/2 - 3 y = 0 - 3 y = -3 and (0, -3) is another ordered pair

jimthompson5910 (jim_thompson5910):

you are correct on both

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

just get the values that they gave you for x | | coordinate pairs are ( x , y) | --- V (2,_) (0,_) substitute /(plug it) into the equation wherever you see "x" y= 1/2 x - 3

OpenStudy (anonymous):

you should then see your answer